9. If the potential at a point which is 1m from a charge is 1 volt, then the potential at a point which is 2m from the same charge will be:


1
Expert's answer
2021-09-12T07:24:49-0400

The electric potential due to a point charge is given by;

"V=\frac{kQ}{r}" Where k =9.0×10^9 Nm²/C²

From which "Q=\frac{Vr}{k}"

Given for 1m=1volt,

="\frac{1V*1m}{9.0×10^9 Nm²/C²}"

=1.1111*10-10C

Now for r=2m?

"V=\frac{kQ}{r}"

"=\frac{9.0×10^9 Nm²/C²*1.1111*10^{-10}C}{2m}"

=0.5volts